OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). sin 2x - sin 4x = 0

OpenStudy (anonymous):

I'm up to sin2x=2sin2xcos2x

OpenStudy (dumbcow):

good, now factor out the sin2x --> sin2x (1 - 2cos2x) = 0

OpenStudy (dumbcow):

sin2x = 0 ----> 2x = 0,pi, x = 0,pi/2 1-2cos2x = 0 ...

OpenStudy (anonymous):

How do you know that 2x=0, π

OpenStudy (anonymous):

and that x=0, π/2?

OpenStudy (dumbcow):

sin(angle) = 0 when angle is multiple of pi angle = pi*n 2x = pi*n x = (pi/2)*n

OpenStudy (anonymous):

Okay, that makes sense

OpenStudy (anonymous):

How do I solve for the sin2xcos2x-1=0?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Anyone please?

OpenStudy (dumbcow):

what i already did? where did the sin come from, after factoring you have sin2x = 0 1-2cos2x = 0

OpenStudy (dumbcow):

its simple algebra to isolate the "cos2x"

OpenStudy (anonymous):

Wait I thought when you factored out the sin2x there's 1 sin2x left in the 2sin2x

OpenStudy (anonymous):

After factoring out I got sin2x(sin2xcos2x-1)

OpenStudy (dumbcow):

no that is wrong, if you multiply that out you will have a sin^2

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (anonymous):

I don't see it though